Can you please help me identify the x intercepts of this equation
y=2(x+4)^2+3
I really appreciate any help...I think I have to find the factors, but I don't know where to start when it's in this form. Thank you so much!

Answer by mananth(16946)

y=2(x+4)^2+3
y=2(x^2+8x16)+3
y=2x^2+16x+32+3
y=2x^2+16x+35
to find X intercept plug y=0
2x^2+16x+35=0
solve by quadratic formula
b^2-4ac <0
x values are not real


therefore the graph has no X intercepts.